OpenStudy (anonymous):

help please :( The electric force between a charged foam cup and a positively charged metal sphere is 2.12 newtons. If the charge on the cup is 2.0 × 10-6 coulombs and the charge on the metal sphere is 2.5 × 10-6 coulombs, what is the distance between the two? (k = 9.0 × 109 newton·meter2/coulombs2

OpenStudy (anonymous):

\[F = \frac{q_1q_2}{4\pi\epsilon_0r^2} = k\frac{q_1q_2}{r^2}\] Rearrange for r, q1 and q2 are the charges of the two objects

OpenStudy (anonymous):

thank you helped alot

OpenStudy (isaiah.feynman):

\[r=\sqrt{k\frac{ q_{1}q_{2} }{ F }}\]